OpenStudy (kewlgeek555):

I Need Help With Functions? Algebra I - Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41… Explain whether or not this sequence can be considered a function.

OpenStudy (kewlgeek555):

This is what I put: I think this sequence can be a function. I am not sure how to explain this function, though. Hm. F(n) F(1) = 5 F(2) = f(1) + 1 F(3) = f(2) + 2 F(4) = f(3) + 3 F(5) = f(4) + 4 And so on and so on?

OpenStudy (anonymous):

@ash2326 return and help him. I need to go to bed x_x

OpenStudy (kewlgeek555):

I have a feeling I am wrong.

OpenStudy (anonymous):

the fact that the differences go up one each time is a good clue that this is a quadratic funcion

OpenStudy (kewlgeek555):

Okay, but did I write the distributive formula correctly?

OpenStudy (anonymous):

you have it right but you want to be looking for something like \(f(x)=ax^2+bx+c\)

OpenStudy (anonymous):

oooh i see, i didn't read the question carefully you do not have to actually find the function do you? only say that it is one

OpenStudy (anonymous):

in which case you are good with your answer it is a function because you came up with a recursive rule for it

OpenStudy (kewlgeek555):

Thank god. Thank you Mr. @satellite73

OpenStudy (anonymous):

yw